Volume 2, Number 1 February 1993

TEX and TUG NEWS

for and by the TEX community

A Publication of the TEX Users Group Electronic version TEX and TUG NEWS

Mission Statement

The TEX Users Group (TUG) provides leadership:

1. to encourage and expand the use of TEX, , and related systems

2. to ensure the integrity and portability of TEX, , and related systems

3. to foster innovation in high-quality electronic document preparation

TEX and TUG NEWS is a newsletter for TEX and L^{A}TEX users alike: a forum for exchanging information, tips and suggestions; a regular means of communicating news items to one another; a place where information about TEX and TUG can be quickly disseminated.

Throughout the newsletter “TEX” is understood to mean TEX, L^{A}TEX, AMS-TEX, and other related programs and macros. TEX and TUG NEWS is produced with the standard L^{A}TEX distribution, and is to be as portable a document as possible.

The entire contents of this newsletter are being placed in the public domain. The source file of this issue will be placed in the aston, shsu, and stuttgart archives, as well as at the heidelberg, labrea, and ymir archives. Copying and reprinting are encouraged; however, an acknowledgement specifying TEX and TUG NEWS as the source would be appreciated.

Submissions to TEX and TUG NEWS should be short, the macros must work, and the files must run without special font or graphics requirements: this is to be a portable newsletter (the new font selection scheme has not yet been implemented). Correspondence may be sent via e-mail to tug@math.ams.org with the subject line NEWSLETTER. Regular mail should be addressed to the Publications Committee, c/o TUG Office at the address below.

TEX Users Group  Phone: 805-963-1338
P.O. Box 869 FAX: 805-963-8358
Santa Barbara, CA 93102 USA E-mail: tug@math.ams.org

TEX is a trademark of the American Mathematical Society.

TEX and TUG NEWS (ISSN 1065-240X) is published four times a year, February, April, July and October, by the TEX Users Group, 735 State Street, P.O. Box 869, Santa Barbara, CA 93102, U.S.A. The 1993 dues for individual members are $60.00 for regular members and $30.00 for students.

Second-class postage paid at Santa Barbara, CA, and additional mailing offices. Postmaster: Send address changes to the TEX Users Group, P.O. Box 869, Santa Barbara, CA 93102, U.S.A.

Membership in the TEX Users Group is for the calendar year, and includes all issues of TUGboat and TEX and TUG NEWS for the year in which membership begins or is renewed. Individual membership is open only to named individuals, and carries with it such rights and responsibilities as voting in the annual election.

Editorial

Welcome to 1993! This is the beginning of TTN’s second year of existence, and there will be a few changes as the issues unfold.

Now that I have the added duties of president of the user group, I will be weaning myself away from TTN so that by next year there will be a new editor in place here. If you feel you can produce 4 issues of a 32-page newsletter, and do it on time (January, April, July and October are our target dates), please send your particulars to me.

Other changes will probably include some shuffling of our columns and adding a few new columnists. Over the past month or so, I’ve received some very thoughtful and constructive comments and suggestions on how to make TTN more useful. Most common complaint is that there’s still not enough basic information (examples, problems and solutions, descriptions) on how to do this or that in (L^{A})TEX.

One way to address this is to get a regular columnist for “Hey — it works!” to take care of submissions sent by users. As well, we plan to add a new column, rather more like a question and answer, or frequently requested solutions. We hope to have people in these two slots by the next issue — so watch for them!

And now for an overview of this month’s issue …

Grab your keyboards and go look at what Peter Flynn’s got in his column this month — a nifty little style file which should appeal to those of you who want to spice up your page layout. And Peter Schmitt has more books for you to think about owning.

Claudio Beccari is back, with more on tables in L^{A}TEX. Actually, it’s more of a commentary and follow-up on something which Jackie Damrau made available last year (see TTN 1,2:16).

Part II of Yannis Haralambous’ answers to questions about DC fonts is here (Part I was in the December issue of last year). If you have any questions not asked here, send them in and we’ll find some answers.

Chris Carruthers, who takes care of electronic distribution of TTN and also serves as a mean proofreader, has written an update on TTN’s electronic availability.

Malcolm Clark was off in the Ukraine last fall, at the invitation of the Shevchenko Scientific Society, and wrote up his experiences, both rueful and otherwise. And there’s news from GUTenberg, with a report from their new interim president, Alain Cousquer.

Speaking of GUTenberg … I was in error in my last editorial about GUTenberg’s involvement with a spring meeting on electronic documents and the law. It was Bernard Gaulle who pointed out that the meeting was actually being sponsored by AFCET. Curious coincidence that halfway around the world, on Feb. 1–2, in Toronto, Canada, a major conference on a very similar theme has been scheduled: “The Ethics of Scholarly Publishing: A Symposium”. With the very plastic nature of electronic information exchange, it is very easy to stray into murky and confusing waters. Interesting topic …

We also have an update on what’s been going on in Russia with CyrTUG: Irina Makhovaya has sent us a report on their latest meeting, and their plans for 1993.

And news about TUG ’93 is starting to be distributed. We have an update on pp. 20–21 — and also check Cover 4.

On the newsletter front, Jackie Damrau was recently awarded one of four DECUS Board of Directors’ Gold Medals of Honor for her volunteer efforts during that organisation’s meeting in Las Vegas last December. Volunteerism often seems to offer more intangible than tangible rewards — recognition is therefore doubly sweet. Congratulations, Jackie!

And with the new year we see a few changes in familiar faces. Malcolm Clark has now left the board, but is busy with the TUG ’93 conference committee and preparations for the meeting at Aston this July. And Malcolm is also preparing a book — on L^{A}TEX — so there’s no doubt he’ll keep on top of things in the TEX community.

Bernard Gaulle, founder of the French-speaking TEX Users Group GUTenberg, is taking a sabbatical from his duties there. However, he assures me he will continue to be active, and I expect we’ll be hearing from Bernard in the future. In the meanwhile, Alain Cousquer has taken over as interim president, and as GUTenberg’s representative on TUG’s board, until elections this summer.

To round out the issue, I thought it might be useful to have a quick reference list to user groups around the world. The information is that which I found in TTN and in the 1991 Resource Directory. If there are any inaccuracies or any new user groups, please contact me.

Enjoy the issue. And think seriously about attending the 1993 TUG meeting at Aston University this coming July 26–30. It’ll be a great meeting!

Christina Thiele

Editor, TEX and TUG NEWS

Did you know …?

… that the very first meeting of TEX users was in February of 1980, in Palo Alto, California. About 50 people attended that meeting; speakers included Don Knuth, Luis Trabb Pardo, Bob McClure, David Fuchs, Richard Zippel, Mike Spivak, and Richard Palais. The first steering committee was elected at that time as well.

Typographer’s Inn

Peter Flynn

University College Cork

                                          cbts8001@iruccvax.ucc.ie

The December/January break gave me the opportunity to reply to a backlog of accumulated mail: thank you to everyone who replied or commented on recent columns, and my apologies if I have missed any of you out.

Punctuation and quotes

I’ve had more mail on this than anything else: the majority of you seem to want it to be treated on the basis of context, so I shall continue to set punctuation inside the closing quotes when it is an integral part of the quote, and outside when it is part of the surrounding verbiage. Our editor points out that the MLA (Modern Language Association) insists on commas and fullpoints going inside the quotes — some people also consider Buckingham Palace a fine building. Several of you said that Hart’s Rules suggest the use of context as a guide, and Bill Woolf says the AMS (American Math Society) recommendations follow this principle, but admits it might lead to sentences like The editor said: “Did he really say ‘Punctuate this way!’ ? ”.

From the mailbox

Christina also asks if newspapers using computers with ‘a’-hyphenation programs count amongst DTP horrors. I’m not sure it’s restricted to the letter ‘a’ by any means, but I would welcome uniquely gross examples such as bathroom which do seem to occur in news work far too often.

Don Hosek at Claremont points out that some of the Usenet newsgroups cover aspects of typography (comp.text.desktop and comp.fonts). I browsed them, but most of the posts seem to be of the “where can I download a copy of Pogmathon Bold Italic for MS-Windows” genre. Traffic on typo-l has been low over the winter, maybe you’ve all been hibernating.

William McKeehan at UTK and several others mention that some academics dislike italics and insist on underlining. \underline{Chacun à son goût}: I suspect they’ve been so used to typewriters they’ve never had the opportunity to change. One advantage of TEX is that it makes it significantly more difficult to do underlining than to do italics.

Glenn Herteg asks about the use of spacing around the em-rule — like this — rather than set solid—like that. I agree with him that it looks better with spacing, but I’ve used a \thinspace in my example rather than a normal wordspace.

If you prefer a wordspace, though, remember to type a like this~--- so that if the rule occurs at a linebreak when justified, the rule is kept on the line with the preceding word (the TEXbook [p. 311 : 12.4]).

Floating free

You can pick up the file flow.sty

 from curia.ucc.ie in pub/tex, or mail
 the request get italic-L log9212 to
 listserv@irlearn.ucd.ie (or contact

me if you have no network access). I’ll buy a pint of the usual at the Aston meeting for the first person to explain how to retain the surrounding paragraph’s sensitivity to \noindent.

The comments people made about flowing text round a dropped initial cap brought some requests for flowing text round boxes against the margin, such as tables, figures and other illustrations. Then out of the blue on the italic-l list came a plea for help on this very subject, so I spent an hour or two trying it out. I think someone has already done this for L^{A}TEX, and I know that

it is much in demand, so my own (very rough) attempt can be picked up as shown. It seems to work in L^{A}TEX as well as plain TEX, but I haven’t tested it exhaustively, perhaps someone could push it to the limits. A brief flick through any magazine will show how extensively this construction is used, but what a pity there is no way to

get TEX to handle an irregular shape automatically for the purposes of flowing text round it. A similar mechanism could be used to handle boxes at the four corners of a page, or spanning n columns of a page at the head or foot, or partway down the page. There was a presentation at the Boston TUG meeting in 1991 about using TEX for page makeup, but floats and flows such as these would be a useful tool for people making magazines and journals.

From the laboratory

I’ve been doing some experimenting with different paper surfaces and typefaces, something which I haven’t done since I had the run of a composing room. What sparked it off was re-reading a comment by Charles Fyffe, “Despite common practice, Times is not suited to art paper” (Fyffe, 1969 : 54). He recommends using these faces on art (glossy) paper: Century, Emerson, Ehrhardt, Goudy Modern, Imprint, Ionic, Poliphilus, Plantin, Plantin Light, Romulus Bold, Egyptian or sans serif (my italics).

I tried out a semi-gloss paper for laserprinters, a sample I was given (so I don’t know the name of it yet). It feeds well, and retains its stability much better than the usual photocopy rubbish. But although not an art paper, it proved Fyffe’s contention because of the reflectivity of the surface. Hairlines tend to break up in electrostatic printing, unlike letterpress or offset litho, partly because of the dot-formation and partly because of the poor surface of most photocopy paper, but this stuff handles them much better. Times (almost anyone’s except Adobe’s) has hair-ended serifs and relatively fine thins, and these lose their apparent definition on glossy surfaces because there is too much reflected light around for your eye to pick them up, even though they may be correctly printed because of the receptivity of the substrate.

I used a 600 dpi HP Laserjet 4 (nice machine, pity about the silly changes to the manual feed, and the appalling paper curling) and Metafoundry Times, which is rather clumsy, so I wasn’t expecting it to be so much of a problem, but sure enough, printing the same text in CM, Times, Stone and Plantin made it quite clear that this Times at least was not suited to very high white or reflective paper. Adobe Times was better, but not much. Hardly surprising, when you consider Times was originally designed for creamy yellow newsprint at sizes no larger than longprimer (10pt) where the squash of ink caused by the letterpress process thickened the thins enough to make it legible. CM performed even worse on glossy paper than Times did, but sans fonts (Helvetica, CM, Gill and Optima) performed excellently.

Mixing it

The CM fonts are based on a design cut in the last century, and used mainly for bookwork, in letterpress, on paper like an esparto or antique wove. The roman still makes a good book face, used for continuous text on an off-white paper at 1200 dpi or better, but most modern DTP work is not like that, and most users want to mix typefaces in one way or another. The sad fact remains that after using TEX for well over a decade, I have yet to find a suitable contrasting face to sort comfortably with it. The closest I have got is Gill Sans Bold for section headings in text work: has anyone else experimented with mixing CM roman with anything other than CM sans?

Elsewhere in Fyffe’s book, he says, “Vincent Steer, a famous typographer of an earlier generation, was once asked, ‘How can I learn to mix type faces?’ and he replied, ‘Learn their histories’. Until you have done so a good rule is to use the related bold of the body face. If there isn’t one, use a sans, but make sure it is bold enough. Few things look worse than a display sans lighter than the body face.” (p. 54). Food for thought.

A challenge

Very few fonts available electronically carry ct or st ligatures, as used in much 18th-century work. With TEX’s precision, and DVIPS’s management of Post-Script’s rotational abilities, would someone like to try faking up ct and st ligatures? (I used the \t “tie-after” accent here.) Of course, extra sorts done in would be the ideal solution: and while we’re at it, has anyone ever fixed CM fonts to have a long s as well?

Bibliography

Fyffe, Charles, Basic Copyfitting. London: Studio Vista, 1969, SBN 289 79705 5.

New Publications

Donald E. Knuth. Literate Programming. (CSLI Lecture Notes 27.) Stanford: Center for the Study of Language and Information, 1992. xiii, 368pp. $19.95 (UK). ISBN 0-937073-80-6 (paper), ISBN 0-937073-81-4 (cloth).

• ‘Literate Programming’ (using the WEB system) was developed by Knuth and used to write TEX and its documentation. This book is a collection of previously published articles and extracts from books, together with additional references and some other new material. In particular, it contains the papers “The Errors of TEX” and the “Error Log for TEX, 1978–91”.

Michael Vulis. Modern TEX and Its Applications. Boca Raton: CRC Publishers, 1992. 275pp., includes diskette. US$32.95 ($39.95 outside the US). ISBN 0-8493-4431-X (paper).

• “A guide to plain TEX (and VTEX), intended for scientists, etc., and support staff who prepare technical texts and documents. It covers all basic topics for understanding the TEX user environment, including fonts and characters, formatting, math mode, macros, terminal and file operations, tables, foreign language capabilities (including Cyrillic), indices, references, tables of contents, scalable typefaces, and graphics. A PC-compatible disk containing examples, extra typefaces, and a ready-to-run restricted [!] version of TEX is included.” [adapted from promotional material]

Antoni Diller. L^{A}TEX Line by Line: Tips and Techniques for Document Processing. (Software Engineering.) Chichester: John Wiley and Sons, 1992. 200pp. US$29.95. ISBN 0-471-93471-2 (softcover).

• “A handbook to the L^{A}TEX text processing system, this easy-to-follow guide teaches all users how to produce a wide variety of documents — from business letters to technical reports. The book offers ready-to-use templates for producing articles, reports, letters and more; explains how to quickly generate tables of contents, indices and bibliographies; and includes many practical examples, an appendix of mathematical symbols and a complete glossary of typesetting terms.” [adapted from promotional material]

G. Grätzer. Math into TEX: A Simplified Introduction Using AMS-L^{A}TEX. Basel: Birkhäuser, 1992. SFr 52.–, DM 57.–. ISBN 3-7643-3637-4. Note: This item was already listed in the last issue. This entry adds the bibliographic data of the co-publisher.

Latest issues of other TEX newsletters

Die TEXnische Komödie: vol. 4, no. 3, November 1992, 56pp (DANTE).

“Hey — it works!”

Experiences with tables

Claudio Beccari

Politecnico di Torino, Italy

                                                 beccari@polito.it

In TEX and TUG NEWS 1,2:16, there appeared a note from Jackie Damrau about handouts on the workshop on tables in L^{A}TEX. I already knew most of that material because I had spent some time in examining the .sty files that came with our implementation of TEX. I found most interesting the discussion on decimal alignment, but I was surprised that in the discussion none of the participants had arrived at an idea I had had, which of course was obvious to my own eyes, but probably would appear as a vicious trick to others.

Jackie Damrau’s handouts contained four solutions to the problem of decimal alignment, and here I quote some of the message that Charles H. Franklin contributed to the workshop:

“From the comments in the replies, it seems I am not the only one who has wanted decimal alignment. So I thought I would post a summary of the various solutions so we can all see what the options appear to be.

1. Use \hphantom to add spaces where needed. E.g.

     \newcommand{\Z}{\hphantom{0}} \newcommand{\M}{\hphantom{$-$}}

then

        {1.0\Z\Z    &   \M0.0003 \atop 0.333      &   $-0.0001$}

line up like you want. […]

2. Use the @{.} argument in the tabular declaration to create two columns with no spaces and a decimal point between them. E.g.

     \begin{tabular}{r@{.}l} 10&7 \\ 1&3 \\ 0&75 \end{tabular}  produces  10.7 1.3 0.75

3. Use DECALIGN.STY […] which implements a d alignment argument for the tabular environment. This d alignment type is in fact a version of (2) above. […] 4. Follow the lead of REVTEX and define \dec x.xx which produces a box of

equal width on both sides of the decimal place. When used with a center aligned column, this produces decimal alignment. […]”

My idea follows the solution outlined above, but is different in some respects:

• the decimal columns are declared as right-aligned instead of centered,

• therefore the only thing you need to add are decimal phantom places on the right only, and you need not care about phantom algebraic signs;

• but the most important point is to provide phantom decimals by means of a single character so that you can maintain a visual alignment also in the ASCII source text.

In order to achieve the third point you need just two local definitions of active characters, as the following:

   \def~{\hphantom{0}}   \catcode`\;=13   \def;{\hphantom{,}}

and use the tilde (~) as a regular digit and the semi-colon as a phantom decimal separator.

The tilde is already an active character so that you need not declare it as such; you just redefine it so as to maintain its definition local.

Tab. 3.10 Coefficienti di conversione fra le unità di misura di TEX e L^{A}TEX. Nelle prime otto colonne i coefficienti sono arrotondati a quattro cifre significative. Nell’ultima colonna appaiono i coefficienti “esatti”, cosicché usando questi ultimi si possono ricalcolare gli altri con tutta la precisione desiderata. Ai fini dei confronti quattro cifre sono più che sufficienti.

       mm      cm      pt     bp      pc       in      dd      cc         sp

mm 1,000  0,100 2,845 2,835 0,2371 0,03937 2,659 0,2216 186 467,98
cm 10,00 1,000 28,45 28,35 2,371 0,3937 26,59 2,216 1 864 679,8
pt 0,3515 0,03515 1,000 0,9963 0,08333 0,01384 0,9346 0,07788  65 536
bp 0,3528 0,03528 1,004 1,000 0,08365 0,01389 0,9381 0,07817  65 781,76
pc 4,218 0,4218 12,00 11,96  1,000 0,1660 11,21 0,9346 786 432
in 25,40 2,540 72,27 72,00 6,023 1,000 67,54 5,628 4 736 286,7
dd 0,3760 0,03760 1,070 1,066 0,08917 0,01481 1,000 0,08333  70 124,086
cc 4,513 0,4513 12,84 12,79  1,070 0,1777 12,00  1,000 841 489,04

Note: this table with its caption is taken from my book “L^{A}TEX: Guida a un sistema di editoria elettronica”. See TTN 1,1: 14.

The choice of the semi-colon requires a little bit of explanation: it is absolutely unimportant to choose the semi-colon or another sign, provided it is assigned the same width as the decimal separator. In non-English speaking/writing countries, the ISO recommendations require the use of the comma as the decimal separator and leave the decimal point only to the English-speaking/writing countries (and to the conventions for real numbers in all programming languages); that is why I defined the semi-colon as a phantom comma. In the States, one might prefer to use the colon as the phantom decimal point:

   \catcode`\:=13   \def:{\hphantom{.}}

In the table example on the previous page, you can see the decimal alignment of the last column, but you also see that there is no separator between the units, the thousands, and the millions; this is due to another ISO recommendation (valid also in the USA, but apparently ignored by US national standards) that formally prohibits the insertion of any kind of separator, except a thin space, between the groups of three digits. If you can decipher the Italian caption, you’ll discover what it is about.

The source code is the following (with some ellipsis in order to keep short lines):

\caption[]{Coefficienti  di  conversione fra le unit\`a
           di misura di \TeX\ e \LaTeX. ...}
\bigskip
\def\C #1 {\multicolumn1{|c}{\tt #1}}
\def\T #1 {\multicolumn1{|c|}{\tt #1}}
\let\t\tt
\def~{\hphantom{0}}  \catcode`\;=13  \def;{\hphantom{,}}
\def\s{\rule{0pt}{2.5ex}}\footnotesize
{\centering\tabcolsep=4.3pt
\begin{tabular}{|l*{9}{|r}|}
\hline
\s      &\C mm & \C cm & ... & \T sp         \\[0.5ex]
\hline
\s\t mm&1,000  &0,100  & ... &   186\,467,98~\\
\t cm  &10,00  &1,000  & ... &1\,864\,679,8~~\\
\t pt  &0,3515 &0,03515& ... &    65\,536;~~~\\
\t bp  &0,3528 &0,03528& ... &    65\,781,76~\\
\t pc  &4,218  &0,4218 & ... &   786\,432;~~~\\
\t in  &25,40  &2,540  & ... &4\,736\,286,7~~\\
\t dd  &0,3760 &0,03760& ... &    70\,124,086\\
\t cc  &4,513  &0,4513 & ... &   841\,489,04~\\[0.5ex]
\hline
\end{tabular}\par}

The little strut \s is used to insert a little space above the letters of every line that is preceded by an \hline.

(L^{A})TEX News

DC fonts — questions and answers (II) _{Yannis Haralambous}

                                        yannis@gat.citilille.fr

This is Part II of a two-part series; Part I appeared in TTN 1,4. The following questions and answers about DC fonts are aimed at providing information from the user’s perspective, and so are intended to be pragmatic rather than theoretical.

Q: I’m using PCTEX (v. 2.93) on an IBM clone. Is there a special installation

routine for the DC fonts? And what about people who use emTEX? A: I’m not a PC specialist, so I may be wrong: there is nothing special about

installing DC fonts. Take the right plain format, create the sizes you need with , put them in the usual places where you used to have CM fonts and run TEX as before.

Q: I’m using a Mac. What about me? A: Euro-OzTEX and DirectTEX include DC fonts. Textures can use DC fonts in

bitmap form. For PostScript DC fonts, contact Blue Sky Research. Q: What about UNIX systems? A: What small computers can do, big ones can do as well. There is nothing

platform-specific with DC fonts, so there is no special version of DC fonts for every platform.

Q: Will I be able to preview my files using DC fonts on my workstation using

the X windows dvi previewer? A: Yes. Unless you have a very old version, which does not support 256-character

fonts (and I’m not even sure if such a version ever existed …) Q: What happens if I send my regular files to someone who’s got the DC fonts

set up — will my file still run there? A: It depends what you mean by ‘files’. If you mean clean TEX or L^{A}TEX source

files using only macros like \rm, \bf, \'e for é (instead of \accent19e), $\Gamma$ for Γ (instead of \char'000), \$ for $ (instead of \string^24), etc., then the files will run and give the expected results. If you used ‘hard’ commands, like \char or \string, then the output may have incorrect characters at those places. If, finally, you mean ‘dvi files using CM-encoded fonts’, then the same fonts will be needed to print or preview. The Technical Council plans to release virtual fonts simulating CM fonts, and based on DC fonts, which will definitely solve this problem.

Q: What if I receive a file which needs the DC fonts and I don’t have them.

What will happen when I run the file? Will it even run? A: As in the previous question, there are three cases:

• you receive the file in ‘clean’ text form,

• you receive the file in ‘not-so-clean’ text form,

• you receive a dvi file using DC fonts.

In cases (a) and (b), same answer as for the previous question. In case (c) the file will not run, period! So, you had better install DC fonts, and then this will never happen …

If you have a text needing fonts you happen not to have, you’re running into a disaster. Of course you will be able to preview by some default font (depending on your driver, usually Times or Courier). What should be avoided is the temptation to make a virtual font to simulate the DC font — virtual fonts should not replace DCs. It’s the other way around: out of DCs we will make virtual CM fonts! So the only advice I can give is: install DCs.

Q: Will L^{A}TEX3 require DC fonts? Will I have to upgrade everything when

L^{A}TEX3 becomes available? Will my old fonts be incompatible? A: I do not have the authority to express myself on L^{A}TEX3-related issues. I

would suggest asking Frank Mittelbach or Rainer Schöpf. As a common mortal L^{A}TEX and NFSS user, I would say that I do not expect L^{A}TEX3 to be dependent upon any encoding. The main common point I see between L^{A}TEX3 and the Cork encoding is that they are very important developments, concerning the whole TEX community, and the survival of TEX more generally.

Q: Do I have to also have the New Font Selection Scheme set up on my machine? Is there a package which has both already set up and easy to install? Is it for IBMs only? Is there one for Macs too? UNIX?

A: NFSS is an independent system for smooth and efficient font management. NFSS can work with every family of fonts, Cork or CM, Greek, IPA, Chinese … As soon as there are orthogonal properties in your font family (size, series, shape, etc.), you can use NFSS, no matter what your fonts are. The NFSS package already provides DC font support.

At the present time, the installation routine for NFSS requires the following changes in order to load DC fonts in the place of Computer Modern:

use preload.dc and fontdef.dc as preload.tex and fontdef.tex; when

prompted for file xxxlfont.sty, reply dclfont.sty. It works!

Concerning IBMs, Macs and UNIX, the answer is simple: NFSS is written in TEX, and this programming language (in which we all trust) is platformindependent so there is no reason why all this shouldn’t run everywhere. (I know — perhaps suspicious readers will say that NFSS involves file names, namely those of the fonts, and these are not always platform-independent.

Nevertheless, all DC fonts have names with fewer than 8 characters, so even DOS can read them — and portability is preserved.)

Q: If I use the NFSS, will I have memory problems?

A: No. NFSS just lets you specify the fonts you use, in a more natural way. For example, if you ask for \bf (boldface) and then \sf (sans-serif) you will get a sans-serif bold font (no need to specify \bf again). If you ask for a combination which is not provided (such as “small capitals sans serif boldface extended italic”) and your fontdef.tex file is well written, you will get a replacement for it. I don’t think this will make you use more fonts than before; it may even prevent you from changing fonts all the time (remember that good old Claude Garamond had only roman, italic and small caps in his repertoire).

Q: How do virtual fonts fit into this DC font scheme of things?

A: Virtual fonts are a concept; DC font scheme is an encoding. I don’t see any direct dependence relation between those two. Virtual fonts are and will continue to be used to re-encode external fonts; for example, PostScript ones. Another use for virtual fonts is to construct characters with diacritics which are not covered by the DC font scheme (this is the case for Welsh, Baltic languages, Esperanto), or to satisfy special typographical needs such as the automatic use of the Dutch ‘ij’ letter. Cork-encoded fonts provide a rich base for virtual fonts. There are other fonts which are not virtual but will be used to create virtual fonts: Cyrillic, African (by Jörg Knappen), IPA, etc.

Q: I’m at a large university where there isn’t all that much TEX support, at least in terms of installing new stuff. Can I do this myself?

A: Of course you can do the installation! Doing a change is always easy; what is difficult is facing the consequences. If your co-users are writing a ‘clean’ code and keep their files in text form, they may not even notice the difference. Otherwise, you should perhaps give them some advice, and perhaps let them make some harmless changes in their source files (‘write a Σ as $\Sigma$, cf. the TEXbook, p. 434, and not in the form of \char"06’).

dvi files should be recompiled, just to be sure. And in any case,

• try the whole thing out first on a personal computer, with 2 or 3 of your colleagues, or

• make a thorough study of your system administrator’s ego; it might be that he will not appreciate such a change, however successful it is …

Q: Will I need any special kind of software in order to get output using these fonts?

A: No. Unless ‘special’ means ‘less than 3–4 years old’.

TTN electronic distribution _{Chris Carruthers}

University of Ottawa

                                            cjc@acadvm1.uottawa.ca

Previous issues of TEX and TUG NEWS are available for anonymous ftp from the following archives: Stuttgart server at rusmv1.rus.uni-stuttgart.de (129.69.1.12) as soft/tex/digests/ttn/ttn1n4.*; from Sam Houston State University, niord.shsu.edu (192.92.115.8), in directory ttn, ttn1n4.* (in VMSSpeak [FILESERV.TTN]TTN1N4.*); and from the Aston server at tex.ac.uk

(134.151.40.18) as [TEX-ARCHIVE.DIGESTS.TTN]TTN1N4.*.

All previous issues have been slightly modified and renamed to make it easier to retrieve them with wildcards (e.g., to retrieve all of volume 1, type: mget ttn1*). The convention is that all names start with the letters ttn followed by a number representing the volume followed by the letter n (‘n’ for number) followed by the issue number, followed by a ‘.’, followed by an extension — usually tex or sty. This naming convention also corresponds with Knuth’s suggestion in the TEXbook (p. 278 in my copy). All subsequent issues, including this one, will follow this convention.

Note: As of this writing you will not be able use mget ttn* at niord without specifying a volume number, because of other files found in the ttn directory.

Files to print up font tables _{Christina Thiele}

                                           cthiele@ccs.carleton.ca

In the last issue of TTN (p. 14), I mentioned that having a file to print up all the characters in a given font was a useful little tool. I’ve had a number of requests for the TEX-able version I have. And now we’ve had an offer of two versions for L^{A}TEX users, from Claudio Beccari:

FONTTAB.TEX and EFONTTAB.TEX run under L^{A}TEX and are intended to produce the font tables for 128 character fonts (the former one) and for the extended 256 character fonts (the latter one). Both are interactive so that the user can specify a different font name at each prompt or the word “stop” (without quotation marks) when he/she is through.

English instructions are in the preamble of each file. Don’t worry about leftover commands that are useless — I did not want to spend too much time in cleaning up a couple of working sets of commands, although I am sure that one can do much better than that.

I hope these files will be useful to you.

If you would like one or more of these files, send me a message.

News from Around …

uk ⟶ ue Malcolm Clark

                                               malcolmc@wmin.ac.uk

One of the delights of having your name crop up in TUGboat or other TEXrelated publications is that you find visitors turning up on your doorstep. Sometimes this is a metaphorical doorstep, when e-mail from far away places appears in your mailbox, but from time to time real people drop by. Late last year, Yuri Melnichuk from Lviv^{1} in the Ukraine turned up. As a result of this contact I visited Lviv at the end of September last year, at the invitation of the Shevchenko Scientific Society. The Society had organised one of its series of conferences on the Ukrainization of Computers.

I was surprised how many of my preconceptions were invalid. On the other hand, I had no very clear idea of what to expect. The history of the Ukraine is not widely taught in British schools. It will not have escaped your notice that the former Soviet Union is no more: it is often unclear just what relationship there is among the constituent units of the former union, or even where the boundaries of the units are. The Ukraine is bounded by Poland (which once included parts of western Ukraine), Romania, Slovakia (assuming this article appears after January 1st, 1993, when Czechoslovakia separates into the Czech lands and Slovakia), Hungary (it was part of the Austro-Hungarian empire too), Moldavia, Byelorussia, Russia itself, and the Black Sea.

In Lviv there is plenty of evidence of the influence of the Austro-Hungarian empire. The trams and the architecture kept reminding me of Vienna, although the Cyrillic signs provided a constant ‘foreign’ undertone: Lviv is well within the world of the Cyrillic alphabet. You may think that all Cyrillic is alike. Far from it: there is a great deal of commonality, but Russian (for example) has some characters which are not used in Ukrainian, and Ukrainian has some of its own characters, too. My simplistic hopes that I could take the CyrTUG Cyrillic TEX distribution which I had obtained the previous week at the EuroTEX conference in Prague, were diluted. In fact, there were even more problems there: I took 3.5′′ high density disks: it took a few days to locate a suitable disk drive. Real floppies — those 5.25′′ disks you have almost forgotten about — are by far the most widely available. A compound problem was that the CyrTUG distribution had been intended for 360 Kbyte floppies and I copied the equivalent of two disks to one (can you guess what happened to non-identical files of the same name?).

I have been slowly working myself into the realisation that the way forward is through Unix and PostScript. Visiting Prague and Lviv reminded me that this is not a universal solution at present. The de facto operating system is MS-DOS, and the common output device is a dot matrix printer. I even have to admit that my disliking of laser printed output as a master for book production was mitigated. Frankly, it is far more important to generate a range of books in Ukrainian (most are in Russian) than to strain at the altar of quality. In a country where my total bill for accommodation amounted to something less than $10 for eight nights, you can probably guess that the public domain nature of TEX was rather attractive, too.

1Lviv is the Ukrainian transliteration; some readers may know of this city as Lvov –Ed.

From my point of view, the most important practical result of the week was when the conference recommended the establishing of a Ukrainian TEX user group. It is quite appropriate that Lviv should be a leader here: Ivan Federov, who had worked in Moscow, settled in Lviv in 1572, and produced a New Testament, using Cyrillic fonts. His role in the history of Cyrillic books has been described as ‘crucial’.

My hosts ensured that I had an excellent time, despite my lack of Ukrainian. I did at least discover the word for beer (almost the same as the Czech word). The major short-term problem is one of transport. E-mail to the Ukraine is a bit erratic (at least, from where I am), and the post is not reliable. If we can overcome this ‘simple’ problem, I think we can look forward to much useful exchange. I have a feeling I will be going back, armed with 5.25′′ disks, more style files, and a phrase book.

I must acknowledge the help and assistance of Ivan Basarab, Yuri Melnichuk, the organisers of the Conference, and the British Council.

News from GUTenberg _{Alain Cousquer}

Université de Lille I

                                                  cousquer@lifl.fr

A number of changes have been occurring inside GUTenberg, the user group for French speakers. Late last year, it was announced that the GUTenberg archives, formerly residing at listserv@dhdurz1 would move to Rennes. It is to be run by volunteers, on an experimental basis: Jacques Beigbeder, Bernard Gaulle, Yannis Haralambous, Michel Lavaud, and Eric Picheral. Its purpose for now is to serve as an archive for distributing software. For anonymous ftp access, the address is ftp.cicb.fr (IP address: 129.20.128.2); directory is pub/GUTenberg (or pub/gut). The “README” file is called “ALIRE”.

At the end of December, the editor of the Cahiers GUTenberg announced that, due to heavy workloads, no issue of the Cahiers had appeared since no. 13, in June 1992. The 14th issue of the Cahiers will be papers presented at the EuroTEX conference held last September in Prague. The 15th issue will appear in early 1993. Starting with no. 16, the Cahiers will appear as thematic issues. Topics currently in preparation include: “ligatures and contextual characters” (in collaboration with the Didot project), and on “TEX and Arabic characters”.

As well, a new bimonthly publication, La Lettre de GUTenberg, is being prepared. It will cover news about the association and related issues, and will be sent free to all paid-up members of GUTenberg. The first issue is slated for January 1993.

And finally, Bernard Gaulle, who began the user group as an informal group in 1984, and has served as GUTenberg’s president since its formal creation in 1988, has taken a sabbatical leave, and handed in his resignation as president, as of 31 December 1992. In 1989, he worked to organise a European TEX meeting in Paris, and later that year, joined TUG’s board at Stanford, along with several other heads of European user groups. But while he leaves both GUTenberg, and TUG’s board, he will certainly not be leaving the use or users of (L^{A})TEX. Alain Cousquer (University of Lille I; cousquer@lifl.fr) will serve as interim president and as GUTenberg’s representative on TUG’s board.

Reports on Meetings

BoF session at TUG’92 _{Paula Gudder}

Denver, Colorado

Last July, at the Portland meeting of the TEX Users Group, about 20 people involved in TEX consulting held a BoF (“Birds-of-a-Feather”) session. The group included those who are now consultants and those who were considering starting some type of small business. Each participant gave a brief background of their situation. Then various concerns were listed for discussion. These included: how to charge; how to find work; how to network within the TEX community and other related communities; how to keep from working in a vacuum.

A sheet with the business cards and/or information of each participant was printed and made available after the meeting to all conference members. For information, contact Arthur Ogawa: ogawa@orion.arc.nasa.gov; 415-691-1126; FAX: 415-962-1969.

NTG meeting: 19 Nov. 1992 Meppel, The Netherlands

Kees van der Laan

                                                 cgl@rugr86.rug.nl

Prior to the meeting, the typographic museum at Meppel was visited — a report about that visit alone would be worthwhile. Demonstrations of the old typographic crafts and machines of the late 19th century and early 20th century were enjoyed. Beautiful! The museum employees themselves had worked with those machines in their younger days. Those who happen to visit Holland some day should certainly pay a visit to that museum, a real must! And when your lucky day is 22nd December, then the place is all lit by candlelight. Can you imagine a cosier walk down history’s lane?

After lunch, the main dish was served by Frank Mittelbach and Yannis Haralambous. Frank elaborated upon the L^{A}TEX3 project and his research in computer-assisted typography. He certainly spotted some essential weaknesses in computer-assisted typesetting. Yannis rehearsed on virtual fonts and captivated most of us with his ScholarTEX. We are all longing for the write-ups to be included in MAPS 93.1.

Between mouthfuls, some business was also done. The budget was agreed upon, the volunteer stuff channelled, and the minutes approved. Along with tea, the usual small talk took place, making us all laugh and feel happy.

Some 60 members attended. Theo Jurriens did a great job in organizing this happening. The hosts — Boom Pers — were great in their modest and kindly offered, but nonetheless much appreciated, hospitality. Our secretary, Gerard van Nes, was interviewed by the local newspaper. You know the sort of thing: about TEX, typesetting, the universe and all those important issues typesetters dream of. For the die-hards the day ended with the traditional meal, in one of the Chinese restaurants Meppel can be proud of, giving everyone the opportunity to speak about making beer, wine or whatever amateur typographers do besides typesetting with (L^{A})TEX.

The 1993 meetings are scheduled for: 10 June at De Bilt, hosted by KNMI, with the theme “From Font to Book”; and 18 November at Den Bos, hosted by OCÉ, where the theme will be “The (L^{A})TEX Working Environment”.

News about CyrTUG and Russian TEX users

Irina Makhovaya

CyrTUG Executive Director

CyrTUG (Cyrillic TEX Users Group; in Russian: Associaciia Pol’zovateleĭ Kirillicheskogo TEX’a) was born in May 1991. Now there are about 50 individual and 14 institutional members; among the latter: Mech-Math Faculty of Moscow State University, Keldysh Institute (Moscow), Institute of Mathematics (Siberian branch, Novosibirsk), Ioffe Physics & Technical Institute (St. Petersburg), Phys-MathLit, Nauka Publishers (Moscow), and Mir Publishers (Moscow).

CyrTUG’s President is Professor Joseph V. Romanovskiĭ, mathematician (St. Petersburg University) jvr@or.math.lgu.spb.su; Executive Director is Irina A. Makhovaya, mathematician (Mir Publishers, Moscow) irina@mir.msk.su.

CyrTUG is a non-profit social community of Russian TEX users. According to CyrTUG statutes, our main purposes are to arrange courses, seminars, meetings, etc.; and to publish textbooks and other TEX materials for non-profit purposes.

18 TEX and TUG NEWS Vol. 2, No. 1, 1993

CyrTUG held its annual meeting in Moscow at the Central Economics & Mathematics Institute (CEMI) on October 20–22, 1992. The meeting was attended by 53 members of the group. In his opening speech, the President, Joseph Romanovskiĭ, noted the growing interest in TEX in Russia and other C.I.S. countries, and expressed the hope that the CyrTUG-93 meeting would gather an even larger audience. Irina Makhovaya, Executive Director, reported on the work of the Board of Directors, and Michael Vinogradov reported on the work of the auditing commission. Both reports were unanimously approved by the meeting.

In addition to the business meeting, the following papers were delivered:

– A.B. Khodulev (Keldysh Institute, Moscow) The public domain Cyrillic TEX.

– M.S. Doubson (CEMI, Moscow) The EuroTEX-92 conference.

– S.A. Strelkov (Keldysh Institute, Moscow) Some aspects of TEX Russification.

– M.M. Vinogradov (Institute of Economics and Forecasting, Moscow) Eight-bit fonts and a user shell for IBM PC.

– E.V. Pankrat’ev (Moscow University) Publishing with TEX in the Math Department of Moscow State University.

– V.A. Rozov (Space Research Institute, Moscow) Electronic archives.

– M.S. Doubson (CEMI, Moscow) Cyrillic fonts in PostScript.

– E.M. Yankovskiĭ (Mir Publishers, Moscow) Using TEX in translating science and mathematics from Russian into English.

– V.M. Rudenko (IPM, Moscow) Working with TEX at the Institute of Problems of Mechanics.

– I.A. Makhovaya (Mir Publishers, Moscow) Publishing with TEX at Mir Publishers.

– J.V. Romanovskiĭ (St. Petersburg University) TEX and e-mail.

– V.G. Perepelkin (Institute of Mathematics, Novosibirsk) The Siberian branch of CyrTUG, SibTUG.

– K.E. Pankrat’ev (Moscow University) A device for optical character recognition.

– E.Yu. Khodan (PhysMathLit, Moscow) Future plans for publishing materials on

TEX.

The meeting agreed on the following plan for 1993:

• To prepare a series of booklets (may be in “soft” variant) for beginners using the TEX, AMS-TEX, and L^{A}TEX packages, especially for the Cyrillic versions.

• To organize several CyrTUG courses for beginners and advanced TEX users with instructors from within and outside Russia.

• To arrange additional services for TEX users who have access to e-mail, such as support of a TEX archive and a Cyrillic FAQ.

• To arrange for regular distribution (at least four times a year) through e-mail or the postal service of an information bulletin for members of CyrTUG.

• To give the right to members to receive xerox copies (free of charge) of some pages of printed materials in the possession of the Board of Directors.

• To hold a conference with the participation of foreign TEXperts.

CyrTUG also informs its regular and potential members that, to become a full member of CyrTUG, the following entrance and annual fees must be paid through a savings bank or the postal service:

entrance annual
for individual members (regular)  800 rb  500 rb
for full-time students  400 rb  250 rb

for institutional membership:

– scientific and educational institutions 17000 rb 10000 rb
– other organisations  42000 rb 25000 rb

For foreign citizens (outside C.I.S.), the entrance fee is set at DM50.– (cash). This fee will be spent completely for the support of Russian specialists developing public domain TEXware. At present, there are two foreign citizens now members of CyrTUG.

On presenting a copy of the payment check or money order and 20 360KB diskettes (for an individual member) or 40 360KB diskettes (for an institutional member), you will be supplied with the CyrTUG version of the TEX package with Latin (English) and Cyrillic fonts (public domain) and the right to receive on-line support and necessary information. The difference in the number of diskettes is due to the fact that individual members are supplied with fonts only for dotmatrix printers while institutions are supplied, in addition, with fonts for laser printers. Institutional members are also supplied with M. Spivak’s book “The Joy of TEX” in English, free of charge. Each institute has the right to send 10 representatives to CyrTUG meetings.

For further information, contact:

Executive Director:

Irina Makhovaya, CyrTUG  irina@mir.msk.su
Mir Publishers  Tel: 095 286-0622, 286-1777
2, Pervyĭ Rizhskiĭ Pereulok,  FAX: 095 288-9522

Moscow 129820, Russia

UK-TUG meeting: April 6th – 8th _{Philip Taylor}

Royal Holloway and Bedford New College

                                          P.Taylor@Vax.Rhbnc.Ac.Uk

Originally scheduled for March 29th – 31st in Glasgow, the UK-TUG meeting on “TEX for non-American languages; in theory and practice” will be held a week later, and at Royal Holloway and Bedford New College (‘The Country Campus of the University of London’). The new date is timed to coincide with the beginning of Spring, when the campus is at its very finest.

The programme is the same as previously announced; April 6th is reserved for travelling, registration and the conference dinner; April 7th will consist of a series of talks (with plenty of time for discussion) from an internationally famous panel of invited speakers, including Dominik Wujastyk (who, together with Graham Toal, is responsible for the authoritative U.K. hyphenation patterns, and who is also a Sanskrit scholar), Bernard Gaulle (until recently president of GUTenberg, the French-speaking TEX Users’ Group, and an authority on good French typographic practice), and Yannis Haralambous (who is the author of ScholarTEX, and a renowned authority on ).

On April 8th, two concurrent tutorials will take place, one on “ in Theory”, led by Yannis, and one on “ in Practice”. The former will be concerned with the design and implementation of new fonts through the medium of , whilst the second will be more practical and aimed at answering questions such as “How do I generate all the Computer Modern fonts at 600dpi for my new laser printer?”.

Royal Holloway and Bedford New College is situated conveniently close to London Heathrow Airport, and is also served by regular fast trains from London Waterloo (circa 30 minutes). Accommodation will be en-suite, and all meals will be included in the price.

For further information, contact Philip Taylor at:

The Computer Centre, RHBNC  Tel: +44 784 443172
University of London  Fax: +44 784 434348

Egham Hill, Egham

      Surrey TW20 0EX  P.Taylor@Vax.Rhbnc.Ac.Uk

United Kingdom

TUG ’93: 26th – 30th July Birmingham, U.K.

World-Wide Window on TEX

The 14th Annual TEX Users Group Meeting is coming soon! Aston University in Birmingham (United Kingdom) will be the venue for the 1993 TUG conference. Aston is the home of the ‘Aston Archive’, one of the largest collections of electronic TEX miscellanea in the known universe.

The world-wide aspects of TEX, L^{A}TEX and are underlined by holding the conference at one of the main centres of the electronic web outside of North America, whilst windowing systems in various forms allow us to exploit more visually oriented methods of employing the TEX tools.

It is also hoped that there will be a contribution to the conference from the Didot project, further extending the range of topics to include digital typography and font creation.

The conference will feature the normal paper presentations, whilst workshops, poster displays, courses, panels and ‘birds of a feather’ sessions will also form integral components of the main conference.

There will be a selection of TEX Users Group courses organised in conjunction with the meeting. It is intended that courses will run both during the week preceding the conference and during the week following; the courses offered and their timetabling are not yet fixed, but we would hope at the very least to be able to offer courses (both intensive-introductory and advanced) on TEX, , and L^{A}TEX; other courses (e.g. document design) are still being discussed.

If you have a paper which you would like to present, please contact the programme committee urgently at TUG93-Proposals@Vax.Rhbnc.Ac.Uk (see outside back cover for snail-mail address). Deadline for paper proposals is … was … February 26th, 1993.

TUG ’93 will take place on the Aston University campus in central Birmingham, with excellent travel connections. Birmingham is situated very conveniently for access to all parts of England, especially the Midlands areas including Stratford and Oxford, as well as much of Wales. Participants will be given details of local events and sightseeing suggestions, including a guide to the culinary treats offered by Birmingham’s multi-ethnic society.

Social events will include dinners, receptions, and a mystery trip to an unexpected face of Birmingham. Accompanying persons will be helped to have a rewarding time in the area, and will also be offered a beginners’ course in TEX; this course is available as well to those who would like a crash course in TEX prior to the conference proper.

The approximate cost of attendance at the conference itself will be $100, or $250 including accommodation and all meals apart from evening dinners. The normal accommodation provided will be on campus in the University student residences, but alternative arrangements (e.g. for executive accommodation on campus, or hotel accommodation) are also possible; full details of these and other options are given on the booking form which is available on request.

Full computer facilities, including electronic mail and network connections, will be available to participants. The rich holdings of the UK TEX Archive will, of course, be easily accessible.

The Conference Committee:

Chairman:  Peter Abbott
Administration:  Maureen Campbell
Programme:  Chris Rowley and Malcolm Clark
Editors:  Sebastian Rahtz and Mimi Burbank
Courses organiser:  Carol Hewlett
Publicity and demonstrations: Philip Taylor
Social programme:  David Osborne

Philip Taylor TUG ’93 Committee

4 conferences to be held jointly: 11–15 April 1994

There will be a major gathering of people interested and involved in the burgeoning field of electronic document processing next year, 11–15 April, 1994, Darmstadt, Germany. The four international conferences are:

EP94  Electronic Publishing, Document Manipulation and Typography

           Contact: ep94@gmd.de

RIDT94 Raster Imaging and Digital Typography

           Contact: ridt94@irisa.fr

TEP94 Teaching Electronic Publishing

           Contact: ltsdyson@uk.ac.reading

PODP94 Principles of Document Processing

           Contact: podp94@cs.umd.edu

Conference announcements are already being prepared. The schedule for the EP94 and RIDT94 portions include the following deadlines: submissions must be received by 15 August 1993; decisions by the selection committee will be made known by 15 October 1993; and final version of papers must be submitted by 3 December 1993. Further information may be obtained by sending queries to the addresses provided above.

TUG Board Activities

Christina Thiele

President, TEX Users Group

Around the board table this year are some changes. With my taking on the position of president, we see the departure of Malcolm Clark from the board. Bernard Gaulle, as outlined elsewhere, will be replaced by Alain Cousquer as GUTenberg’s representative on the board.

I first met Malcolm in the person of Cathy Booth — she was acting as his proxy at the 1988 TUG meeting in Montreal, where I first joined the board. It seemed already then that Malcolm was well-known to many people there. I finally met him in person the following year, at TUG’s 10th Anniversary meeting at Stanford. It was a very exciting time that year: the enthusiasm and zeal with which he approached TEX (or rather, τϵχ — first person I met who actually pronounced it that way) was great! And for anyone who’s seen him give papers, that non-stop energy is very inspiring.

While on the board, Malcolm served on a number of committees, including Local User Groups, Membership, Long Range Planning, and Publications. With his board activities now behind him, that enthusiasm is sure to be back in full swing at Aston and elsewhere.

It was also at Stanford where I first met Bernard, along with the representatives of other European user groups: Roswitha Graham of the Nordic group, Kees van der Laan for the Dutch, and Joachim Lammarsch for the Germanspeaking user group. I remember sitting at a board meeting-cum-luncheon, and seeing Bernard with this rather puzzled, yet bemused expression on his face — “so this is what a board is like, with 27 people” or something like that.

In addition to his duties as president of GUTenberg, which he began in 1984, Bernard also found time to be on the Conference Planning committee. I’ve enjoyed his presence, and I will miss his thoughtful participation and direct approach.

Nevertheless, one must move on. These two men have many new projects on their plates. Their contributions to TUG have been considerable, and their continued presence within the TEX community will certainly be followed with interest.

Next Board Meeting

24–25 July, at TUG ’93

Aston University, Birmingham, England

TEX Users Group 1993 Course Schedule

Beginning/Intermediate TEX* Intensive Course in L^{A}TEX*
Boston  April 19–23  San Francisco March 1–5
San Diego  June 14–18  Boston  April 26–30
Boston  August 9–13  Ottawa  August 23–27
Chicago  Oct. 18–22  Boston  Oct. 25–29
Modifying L^{A}TEX Style Files* Adv’d TEX and Macro Writing*
San Diego  June 7–11  San Diego  June 21–25
Boston  Nov. 1–5
TEX for Publishers  Practical SGML and TEX*
San Francisco  March 8  San Francisco March 9–10
Boston  March 22  Boston  Nov. 8–9
New York City June 9
Washington, DC Nov. 12
Book and Document Design  SGML and TEX for
with TEX  Publishers*
Boston Sept. 23–24  New York Nov. 10

*Lab classes — computers will be provided for all students.

• TUG courses are small, with 8–15 students in most classes, and are held at major hotels.

• The dates and locations above are tentative — for more information, contact the TEX Users Group at 805-963-1338 or send a fax to 805-963-8358.

On-Site Courses in TEX and L^{A}TEX from TEX Users Group

• Courses in TEX, L^{A}TEX, SGML and TEX, PostScript, or TEX for Publishers tailored to the needs of your group

• Courses at every level from beginning to advanced

• Five full days of instruction at your site

• One-week course fee includes all instructor fees and expenses plus textbooks and other materials for up to 15 students

• If a properly equipped training facility is not available, TUG will arrange computer rentals and supply TEX or L^{A}TEX software

Upcoming Events

1 March Knuth Scholarship: Deadline for 1993 submissions. (see Cover 2 for address)

Knuth Scholar. Committee c/o TUG Office

                           tug@math.ams.org

9–12 March

DANTE’93:

General Meeting, Chemnitz (near Dresden).

Dr. Wolfgang Riedel

                           wolfgang.riedel@hrz.tu-
                           chemnitz.de

1 April EPodd deadline. Call for papers for special issue of Electronic Publishing: Origination, Dissemination and Design on Active Documents (see TUGboat 13, no. 4, p. 532 for details).

Vincent Quint

INRIA/IMAG

2, rue de Vignate

F-38610 Gieres, France

                           quint@imag.fr
                           Tel: +33 76 63 48 31
                           FAX: +33 76 54 76 15

7–8 April

UKTUG:

“TEX for non-American Languages”. Conference and tutorials, Royal Holloway and Bedford New College (see pp. 19– 20).

Phil Taylor

                           p.taylor@vax.rhbnc.ac.uk

May

UKTUG:

Visit to John Wiley and Sons Ltd., Chichester, England

                           David Murphy d.v.murphy@
                           computer-science.
                           birmingham.ac.uk

10 June

NTG:

11th Annual Meeting: “From Font to Book”. Royal Dutch Meteorological Inst., De Bilt.

Theo Jurriens

                           taj@astro.rug.nl

cont’d … 16–18 June

SSP93:

15th Annual Meeting: “What Will We Be Tomorrow?” Society for Scholarly Publishing Hyatt Crystal City, Washington, D.C.

SSP, 10200 W 44th Ave., #304, Wheat Ridge, CO 80033, USA Tel: 303-422-3914 FAX: 303-422-8894

16–19 June

ACH-ALLC93:

Joint int’l. conference: Assoc. for Computers and the Humanities and the Assoc. for Literary and Linguistic Computing Georgetown University, Washington, D.C.

Dr. Michael Neuman

                           Neuman@Guvax.bitnet or
                           Neuman@Guvax.Georgetown.
                           edu

26–29 July

TUG ’93:

“World Wide Window on TEX” 14th Annual Meeting, Aston University, Birmingham, UK. (details on pp. 20–21; Cover 4)

Chris Rowley, ca rowley

                           @vax.acs.open.ac.uk

Malcolm Clark

                           malcolmc@wmin.ac.uk

15 Aug.

EP94, RIDT94 deadline: Call for papers for major joint conference in Darmstadt, Germany, 11–15 April 1994. (see p. 22 for details).

                           EP94: ep94@gmd.de
                           RIDT94: ridt94@irisa.fr

18 Nov.

NTG:

12th Annual Meeting: “L^{A}(TEX) User Environment”. OCÉ, Den Bosch.

Gerard van Nes

                           vannes@ecn.nl

Note: Also consult the “Calendar” in the previous issue of TUGboat for more dates and details.

TUG Office

New Phone/FAX Numbers

Phone number: 805-963-1338
Fax number:  805-963-8358

Other TEX User Groups

The following information has been culled from previous issues of TTN and from the 1991 Resource Directory (pp. 119–125). Any groups not listed are invited to contact the editor.

CsTUG (Czech)

Jiří Veselý, President Československé sdružení uživatelů TEXu

CsTUG, c/o MÚK UK

Sokolovská 83 CS-186 00 Praha 8, Czechoslovakia

jvesely@cspguk11.bitnet

CyrTUG (Russian)

Irina Makhovaya, Executive Director Associaciia Pol’zovateleĭ Kirillicheskogo TEX’a Mir Publishers 2, Pervyĭ Rizhskiĭ Pereulok Moscow 129820, Russia Tel: 095 286-0622, 286-1777 FAX: 095 288-9522

irina@mir.msk.su

DANTE (German-speaking)

Joachim Lammarsch, President Deutschsprachige Anwendervereinigung TEX e.V. Postfach 10 18 40 BRD-6900 Heidelberg 1 FAX: 06221/56 55 40

dante@dhdurz1.bitnet

Estonian User Group Enn Saar, Tartu Astrophysical Observatory, Toravere EE2444 Estonia

saar@aai.tartu.ew.su

GUST (Polish)

Hanna Kołodziejska, President Polska Grupa Użytkowników Systemu TEX Instytut Badań Systemowych

PAN

ul. Newelska 6

01-447 Warszawa, Poland

                                  Internet: gust@camk.edu.pl
                                  Bitnet: gust@plcamk61

GUTenberg (French-speaking) Alain Cousquer, Interim President Group francophone des Utilisateurs de TEX Association GUTenberg BP 21 F-78354 Jouy en Josas Cedex, France

Nordic TEX Group (Scandinavian Countries) Roswitha Graham KTH (Royal Institute of Technology)

DAB

S-100 44 Stockholm, Sweden

                                  roswitha@admin.kth.se

NTG (Dutch-speaking)

Kees van der Laan, Chair Nederlandstalige TEX Gebruikersgroep Postbus 394 NL-1740 AJ Schagen The Netherlands,

                                  cgl@rug86.rug.nl

SibTUG (Siberian TUG)

Dr. Anatoly Urvantsev, Chairman Computing Center prosp. Lavrent’eva 6 Novosibirsk 630090, Russia Tel: (3832) 350-454 FAX: (3832) 324-259

uran@isi.itfs.nsk.su

TUG (International user group)

Christina Thiele, President TEX Users Group P.O. Box 869 Santa Barbara, CA 93102 USA Tel: ** 805-963-1338 **^{1} FAX: ** 805-963-8358 **

tug@math.ams.org

UK-TUG (United Kingdom) Peter Abbott, Chairman UK TEX Users’ Group Information Services Aston University Aston Triangle Birmingham B4 7ET, England Tel: +44 21 359 5492

                                  p.abbott@aston.ac.uk

1Note: NEW phone and fax numbers.

TEX and TUG NEWS Table of Contents

Editorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Typographer’s Inn Peter Flynn . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

New Publications Peter Schmitt . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

“Hey — it works!”

Experiences with tables Claudio Beccari . . . . . . . . . . . 7

(L^{A})TEX News DC Fonts: Questions and Answers (II) Yannis Haralambous 10 TTN electronic distribution Chris Carruthers . . . . . . . . 13 Files to print up font tables Christina Thiele . . . . . . . . 13

News from Around … uk ⟶ ue Malcolm Clark . . . . . . . . . . . . . . . . . . . 14 News from GUTenberg Alain Cousquer . . . . . . . . . . . 15

Reports on Meetings BoF session at TUG’92 Paula Gudder . . . . . . . . . . . . 16 NTG meeting: 19 Nov. 1991 Kees van der Laan . . . . . . . 16 News about CyrTUG and Russian TEX users Irina Makhovaya 17 UK-TUG meeting: April 6th–8th Philip Taylor . . . . . . . 19 TUG ’93: 26th–30th July Philip Taylor . . . . . . . . . . . . 20 4 conferences to be held jointly: 11–15 April 1994 . . . . . . . 22

TUG Board Activities . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

TEX Users Group 1993 Course Schedule . . . . . . . . . . . . . . . . . 24

Upcoming Events . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

Other TEX User Groups . . . . . . . . . . . . . . . . . . . . . . . . . . 27

Volume 2, No. 1, 1993 TUG ’93

World Wide Window on TEX

14th Annual TEX Users Group Meeting

Aston University, Birmingham, UK

July 26th – 30th, 1993

Aston University in Birmingham (United Kingdom) will be the venue for the 1993 TUG conference. The world-wide aspects of TEX, L^{A}TEX and are underlined by holding the conference at one of the main centres of the electronic web outside of North America, whilst windowing systems in various forms allow us to exploit more visually oriented methods of employing the TEX tools.

The Aston University campus is in central Birmingham, which has excellent travel connections. Birmingham is situated very conveniently for access to all parts of England, especially the Midlands areas including Stratford and Oxford, as well as much of Wales. Social events will include dinners, receptions, and a mystery trip to an unexpected face of Birmingham. Participants will be given details of local events and sightseeing suggestions, including a guide to the culinary treats offered by Birmingham’s multi-ethnic society.

The conference will feature the normal paper presentations, whilst workshops, poster displays, courses, panels and ‘birds of a feather’ sessions will also form integral components. There will be a selection of TEX Users Group courses organised in conjunction with the meeting, which will take place during the week preceding and the week following the conference proper.

Requests for further information (full details of costs, accommodation options, conference registration forms, etc.), as well as requests for space for demonstrations and displays, should all be sent to the following address, as should completed forms and any other written correspondence:

Peter Abbott  phone: +44 21 359 5492
Information Services fax:  +44 21 359 6158

Aston University

        Aston Triangle  TUG93-Enquiries@Vax.Rhbnc.Ac.Uk

Birmingham B4 7ET United Kingdom

Full computer facilities, including electronic mail and network connections, will be available to participants. The rich holdings of the UK TEX Archive will, of course, be easily accessible.

Important details inside: Read more about TUG ’93 in this issue (p. 20–21).